OpenStudy (flcon04):

(x-1) (x^3+ x^2 + x)= ?

OpenStudy (mayankdevnani):

\[\large \bf (x-1)(x^3+x^2+x)=x^4+x^3+x^2-x^3-x^2-x\] \[\large \bf x^4-x=\color{red}{Answer}\]